OpenStudy (anonymous):

Challenge of the day: Give a closed-form expression for the following integral: \[ \int_0^{\frac{\pi}{2}}\frac{\sin^n(x)}{\cos^n(x)+\sin^n(x)}\;dx \] For all \(n\in\mathbb{Z}\) And, of course, prove it. Nicest solution gets a medal and large props. I will post my own solution in 48 hours.

OpenStudy (anonymous):

No attempts? :(

OpenStudy (anonymous):

Last bump for this question, and I'll post up the answer.

OpenStudy (kainui):

I just saw it now, maybe if you had posted this before the weekend I would've attempted, but if it's too late, then it's too late. I don't have a pen with me and I'm at the library, so not going to give a response today that's for sure.

OpenStudy (anonymous):

I can postpone it, if you wish, so long as there's an attempt. I did post it two days ago, though. @Kainui

OpenStudy (kainui):

Haha, well I want to try it, but I doubt I'll remember, I have a busy weekend ahead of me I'm afraid. Just post the answer and I'll be here to enjoy it at least haha.

OpenStudy (anonymous):

That works. @Kainui

OpenStudy (anonymous):

The most interesting part about this integral is that it's near-impossible to evaluate directly, but can be evaluated under inspection. So, we generalize. Let: \[ f(x)=\sin^n(x) \]And \[ g(x)=\cos^n(x) \]And, let: \[ D=\frac{\pi}{2} \]Note that we have the property that, in general \(f(x)=g(D-x)\). So, we now have: \[ \int_0^D\frac{f(x)}{g(x)+f(x)}\;dx\equiv S \]Now, note the integral: \[ \int_0^D \frac{g(x)}{f(x)+g(x)}\;dx \]Under u-substitution, we have: \[ \begin{align} u&=D-x\\ du&=-dx \end{align} \]Hence our integral becomes: \[ -\int_D^0\frac{g(u)}{f(u)+g(u)}\;du=\int_0^D\frac{g(u)}{f(u)+g(u)}\;du=\int_0^D\frac{f(x)}{g(x)+f(x)}\;dx=S \] So: \[ \int_0^D\frac{f(x)}{g(x)+f(x)}\;dx+\int_0^D\frac{g(u)}{f(u)+g(u)}\;du=2S \]Note that the choice of variable of integration is arbitrary, so we have: \[ \int_0^D\frac{f(x)}{g(x)+f(x)}\;dx+\int_0^D\frac{g(u)}{f(u)+g(u)}\;du=\int_0^D\frac{f(x)+g(x)}{g(x)+f(x)}\;dx= \]\[ \int_0^D1\;dx=D=2S \]So: \[ S\equiv\int_0^D\frac{f(x)}{g(x)+f(x)}\;dx=\frac{D}{2} \]Which is fantastic, because it works for any case with the above properties. Now, we previously defined \(D=\frac{\pi}{2}\), giving us: \[ \int_0^\frac{\pi}{2}\frac{\sin^n(x)}{\cos^n(x)+\sin^n(x)}\;dx=\frac{\pi}{4}, \;\forall n\in \mathbb{Z} \]

OpenStudy (anonymous):

There we go. Sorry, the site went down a moment, for me.